OpenStudy (anonymous):

A lot measures 75 ft by 100 ft. It is all lawn except for the house on the lot. The house measures 30 ft by 50 ft. What is the area of the lawn?

OpenStudy (anonymous):

Hello? :_(

OpenStudy (wolf1728):

ALL Area = 75*100 = 7,500 House Area = 30*50 = 1,500 Lawn Area = 7,500 - 1,500= 6,000 Square Feet

OpenStudy (anonymous):

zank uz!

OpenStudy (wolf1728):

u r welcome